# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v1

@@ C.2 Proof of Theorem 2
$\big{(}R^{\prime}_{k}(t)-\bar{h}_{k}\big{)}^{2}$ $=\frac{2}{t^{2}}\sum_{s<u}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s:)\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u:)\mathbf{e}_{k}$
$+\frac{1}{t^{2}}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u:)\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u:)\mathbf{e}_{k}\enspace.$
Define $v(s):=(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s:)\mathbf{e}_{k}$. Since $|v(s)|\leq 1$, it follows that $|\mathbb{E}[v(s)v(u)]|\leq\mathbb{E}[v(u)]|$.
Remark.
We could provide a tighter bound using that $-\bar{h}_{k}\leq v(s)\leq 1-\bar{h}_{k}$, which implies:
$|\mathbb{E}[v(s)v(u)]|\leq\max(\bar{h}_{k},1-\bar{h}_{k})|\mathbb{E}[v(u)]|.$
Using the previous result, we obtain:
$|\mathbb{E}[(R^{\prime}_{k}(t)-\bar{h}_{k})^{2}]|\leq\frac{1}{t^{2}}\left[\sum_{u=0}^{t-1}(2u+1)|\mathbb{E}[v(u)]|\right].$
Given that $\mathbb{E}[v(u)]=(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}^{u}\mathbf{e}_{k}$, we can further bound $|\mathbb{E}[v(u)]|$ as follows:
$|\mathbb{E}[v(u)]|\leq\|\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n}\|\cdot\lambda_{2}(1)^{u}.$
Additionally, we know that:
$\sum_{u=0}^{t-1}(2u+1)\lambda_{2}(1)^{u}\leq\frac{1+\lambda_{2}(1)}{(1-\lambda_{2}(1))^{2}}.$
Combining these results, we derive:
$|\mathbb{E}[(R^{\prime}_{k}(t)-\bar{h}_{k})^{2}]|\leq\frac{1}{t^{2}}\cdot\frac{1+\lambda_{2}(1)}{(1-\lambda_{2}(1))^{2}}\tilde{\sigma}_{n}(\bar{h}_{k}),$
where $\tilde{\sigma}_{n}(\bar{h}_{k})=\|\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n}\|=\sqrt{n\bar{h}_{k}(1-\bar{h}_{k})}$. From $1-\lambda_{2}(1)=\frac{\lambda_{2}}{|E|}$, we have
${\frac{1+\lambda_{2}(1)}{(1-\lambda_{2}(1))^{2}}}\leq 2\left(\frac{|E|}{\lambda_{2}}\right)^{2}.$
Using $\mathbb{E}[(R^{\prime}_{k}(t)-\bar{h}_{k})^{2}]=\frac{1}{n^{2}}\mathbb{E}[(R_{k}(t)-r_{k})^{2}]$ finishes the proof. ∎
@@ D.1 Proof of Lemma 1
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\frac{\sqrt{2}}{ct}\cdot\frac{c_{n}(r_{k})}{\gamma_{k}(1-2\alpha)}$ (20)
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq\frac{1}{2}$ with $a=\lfloor\alpha n\rfloor+\frac{1}{2}$ and $b=n-\lfloor\alpha n\rfloor+\frac{1}{2}$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $c_{n}(\cdot)$ are as defined in Theorem 2.
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|\geq\frac{\sqrt{2}}{ct}\cdot\frac{c_{n}(r_{k})}{\gamma_{k}},$ (21)
@@ D.2 Proof of Theorem 3
Moreover, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$, we have
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\left(\frac{1}{t}+\frac{2}{ct}+\frac{2\log(t)}{ct-2\log(t)}\right)\cdot\frac{\sqrt{2}}{c(1-2\alpha)}\cdot\|\mathbf{K}\odot\mathbf{X}\|.$
where $\mathbf{K}=\left[\frac{c_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n}$ and $c$, $c_{n}(r_{k})$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
$\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{C_{i}}{t}\cdot|X_{i}|$ (28)
$\leq\frac{1}{t}\left\|\mathbf{C}\odot\mathbf{X}\right\|,$ (29)
where $C_{i}=\frac{\sqrt{2}}{c}\cdot\frac{c_{n}(r_{k})}{\gamma_{k}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-\frac{\log(t)}{c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(t)\right\}$. We obtain for all $t>T^{*}$,
@@
Using Lemma 1, we derive for $s>1$, $\Delta{w}_{k}(s)\leq\frac{C_{k}}{s-1}$. We obtain $\|\mathbf{S}(1)\|\leq\|\mathbf{C}\odot\mathbf{X}\|$ and for $s>1$,
$\|\mathbf{S}(s)\|\leq\sqrt{\sum_{k=1}^{n}\frac{C_{k}^{2}X_{k}^{2}}{(s-1)^{2}}}\leq\frac{1}{s-1}\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\leq e^{-c_{2}(t-T)}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+(t-T)\frac{\lambda_{2}(2)}{T}\|\mathbf{C}\odot\mathbf{X}\|$
$\leq\frac{1}{t}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(t)}{c_{2}t-\log(t)}\lambda_{2}(2)\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{1}{t}\left\|\mathbf{C}\odot\mathbf{X}\right\|+\frac{1}{c_{2}t}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(t)}{c_{2}t-\log(t)}(1-c_{2})\|\mathbf{C}\odot\mathbf{X}\|,$
where $c_{2}=\left(1-\lambda_{2}(2)\right)>0$ with $\lambda_{2}(2)$ the second largest eigenvalue of $\mathbf{W}_{2}$ and the matrix $\mathbf{C}$ is defined as $\mathbf{C}=\left[\frac{\sqrt{2}}{c}\cdot\frac{c_{n}(r_{k})}{\gamma_{k}(1-2\alpha)}\right]_{k=1}^{n}.$ This bound simplifies to
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{1}{t}\left\|\mathbf{C}\odot\mathbf{X}\right\|+\frac{2}{ct}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(t)}{\frac{c}{2}t-\log(t)}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c=\frac{\lambda_{2}}{|E|}=2c_{2}$. Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|=\frac{\sqrt{2}}{c(1-2\alpha)}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{c_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n},$ which completes the proof.
@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
$\frac{1}{n}\max\left(\left\lfloor m+\frac{1}{2}-\frac{K_{n}}{t-T}\right\rfloor,0\right)\leq\varepsilon^{*}_{i}(t)\leq\frac{m}{n}\enspace,$
where $T=\frac{4}{c}\log\left(\frac{n}{\varepsilon\delta}\right)$ denote a propagation time with $c$ being the connectivity constant and $\varepsilon:=\tau/\max_{i}|X_{i}|$ denote a tolerance parameter. Moreover, we define $K_{n}:=\frac{c_{n}(m)}{\left(1-\frac{1}{n}\right)c\delta}$.
@@
$\tilde{\varepsilon}^{*}(t)\geq\frac{\max\left(\left\lfloor m+\frac{1}{2}-\frac{c_{n}(m)}{\delta_{1}ct}\right\rfloor,0\right)}{n},$
where we define $m=\lfloor n\alpha\rfloor$ and $\tilde{c}_{n}(G)$ defined in section 3. As $t\to\infty$, note that $p=\lfloor\frac{1}{2}+m\rfloor=m$, which allows us to recover the breakdown point of the $\alpha$-trimmed mean.
@@
$p_{k}(t)\leq\frac{c_{n}(r_{k})}{(a-r_{k})ct}\leq\frac{c_{n}(m)}{(a-r_{k})ct},$
@@
$p_{k}(t)\leq\frac{c_{n}(m)}{(a-r_{k})ct}\leq\delta.$
and obtain
$p=\max\left(\left\lfloor\frac{1}{2}+m-\frac{c_{n}(m)}{\delta ct}\right\rfloor,0\right).$

### v5

@@ C.2 Proof of Theorem 2
$\big{(}R^{\prime}_{k}(t)-\bar{h}_{k}\big{)}^{2}$ $=\frac{2}{t^{2}}\sum_{s<u}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}$
$+\frac{1}{t^{2}}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\enspace.$
Denoting $\tilde{\mathbf{h}}_{k}=\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n}$, define $v(s):=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s)^{\top}\mathbf{e}_{k}$. We first consider the second term: $\frac{1}{t^{2}}\sum_{u=0}^{t-1}v(u)^{2}$. Note that
$v(u)^{2}=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})\tilde{\mathbf{h}}_{k}\enspace.$
Let $\mathbf{J}=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. Since $\tilde{\mathbf{h}}_{k}^{\top}\mathbf{J}=0$, we can simplify:
$v(u)^{2}=\tilde{\mathbf{h}}_{k}^{\top}\mathbf{W}_{1}(u{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\mathbf{W}_{1}(u{:})\tilde{\mathbf{h}}_{k}\leq\tilde{\mathbf{h}}_{k}^{\top}\mathbf{W}_{1}(u{:})^{\top}\mathbf{W}_{1}(u{:})\tilde{\mathbf{h}}_{k}.$
Since the product of two identical permutation matrices is equal to the identity matrix, we obtain $\mathbf{W}_{1}(u{:})^{\top}\mathbf{W}_{1}(u{:})=\mathbf{I}_{n}$ and conclude that $\mathbb{E}[v(u)^{2}]\leq\|\tilde{\mathbf{h}}_{k}\|^{2}$. Now let’s consider the first term: $\frac{2}{t^{2}}\sum_{s<u}v(s)v(u)$. Note that for $s<u$,
$v(s)v(u)=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:}s+1)\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
Taking expectation conditional on $\mathbf{W}_{1}(s{:})$, we have
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\mathbb{E}\left[\tilde{\mathbf{W}}_{1}(u{:}s+1)\right]\cdot\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
Since $\mathbb{E}[\tilde{\mathbf{W}}_{1}(u{:}s+1)]=\tilde{\mathbf{W}}_{1}^{u-s}$, we obtain the following bound:
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]\leq\lambda_{2}(1)^{u-s}\cdot\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
Using the previous derivation, we get:
$\mathbb{E}[v(s)v(u)]\leq\lambda_{2}(1)^{u-s}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$
Combining the two terms, we obtain:
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{t^{2}}\sum_{s<u}\lambda_{2}(1)^{u-s}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$
Note that:
$\sum_{s<u}\lambda_{2}(1)^{u-s}\leq\sum_{u=0}^{t-1}\sum_{d=1}^{u}\lambda_{2}(1)^{d}\leq\frac{t}{1-\lambda_{2}(1)}\enspace.$
Recalling $c=1-\lambda_{2}(1)$,
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{ct}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\leq\frac{3}{ct}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$
Plugging $\|\tilde{\mathbf{h}}_{k}\|^{2}=\sigma_{n}(\bar{h}_{k})^{2}$ and using $\mathbb{E}[(R^{\prime}_{k}(t)-\bar{h}_{k})^{2}]=\frac{1}{n^{2}}\mathbb{E}[(R_{k}(t)-r_{k})^{2}]$ finish the proof. ∎
@@ D.1 Proof of Lemma 1
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\sqrt{\frac{3}{ct}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}(1-2\alpha)},$ (20)
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq\frac{1}{2}$ with $a=\lfloor\alpha n\rfloor+\frac{1}{2}$ and $b=n-\lfloor\alpha n\rfloor+\frac{1}{2}$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{n}(\cdot)$ are as defined in Theorem 2.
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|)\leq\sqrt{\frac{3}{ct}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}},$ (21)
@@ D.2 Proof of Theorem 3
Moreover, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>\log(t/c)\right\}$, we have
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\left(\frac{3}{\sqrt{ct}}+\frac{\log(t/c)}{(ct-\log(t/c))^{1/2}}\right)\cdot{\frac{\sqrt{3}}{\sqrt{c}(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n}$ and $c$, $\sigma_{n}(r_{k})$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
$\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{C_{i}}{\sqrt{t}}\cdot|X_{i}|$ (28)
$\leq\frac{1}{\sqrt{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|,$ (29)
where $C_{i}=\sqrt{\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{i})}{\gamma_{i}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-\frac{\log(\sqrt{t/c})}{c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(\sqrt{t/c})\right\}$. We obtain for all $t>T^{*}$,
@@
Using Lemma 1, we derive for $s>1$, $\Delta{w}_{k}(s)\leq\frac{C_{k}}{\sqrt{s-1}}$. We obtain $\|\mathbf{S}(1)\|\leq\|\mathbf{C}\odot\mathbf{X}\|$ and for $s>1$,
$\|\mathbf{S}(s)\|\leq\sqrt{\sum_{k=1}^{n}\frac{C_{k}^{2}X_{k}^{2}}{(s-1)}}\leq\frac{1}{\sqrt{s-1}}\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\leq e^{-c_{2}(t-T)}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+(t-T)\frac{\lambda_{2}(2)}{\sqrt{T}}\|\mathbf{C}\odot\mathbf{X}\|$
$\leq\sqrt{\frac{c}{t}}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(\sqrt{t/c})}{({c_{2}t-\log(\sqrt{t/c}))^{1/2}}}\lambda_{2}(2)\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{1}{\sqrt{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|+\frac{2}{\sqrt{ct}}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\frac{1}{2}\log(t/c)}{({\frac{c}{2}t-\frac{1}{2}\log(t/c)})^{1/2}}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c_{2}=\left(1-\lambda_{2}(2)\right)>0$ with $\lambda_{2}(2)$ the second largest eigenvalue of $\mathbf{W}_{2}$ and the matrix $\mathbf{C}$ is defined as $\mathbf{C}=\left[\sqrt{\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}(1-2\alpha)}\right]_{k=1}^{n}.$ This bound simplifies to
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{3}{\sqrt{ct}}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(t/c)}{(ct-\log(t/c))^{1/2}}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c=\frac{\lambda_{2}}{|E|}=2c_{2}$. Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|={\frac{\sqrt{3}}{\sqrt{c}(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n},$ which completes the proof. ∎
@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
$\frac{1}{n}\max\left(\left\lfloor m+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{i}(t)\leq\frac{m}{n}\enspace,$
where $T=\frac{4}{c}\log\left(\frac{n}{\varepsilon\delta}\right)$ denote a propagation time with $c$ being the connectivity constant and $\varepsilon:=\tau/\max_{i}|X_{i}|$ denote a tolerance parameter. Moreover, we define $K(\delta):=\frac{\sqrt{3}\sigma_{n}(m)}{\left(1-\frac{1}{n}\right)\sqrt{c}\delta}$.
@@
$\tilde{\varepsilon}^{*}(t)\geq\frac{\max\left(\left\lfloor m+\frac{1}{2}-\frac{c_{n}(m)}{\delta_{1}\sqrt{ct}}\right\rfloor,0\right)}{n},$
where we define $m=\lfloor n\alpha\rfloor$ and $c_{n}(m)=\sqrt{3}\sigma_{n}(m)$. As $t\to\infty$, note that $p=\lfloor\frac{1}{2}+m\rfloor=m$, which allows us to recover the breakdown point of the $\alpha$-trimmed mean.
@@
$p_{k}(t)\leq\frac{\sqrt{3}\sigma_{n}(r_{k})}{(a-r_{k})\sqrt{ct}}\leq\frac{\sqrt{3}\sigma_{n}(m)}{(a-r_{k})\sqrt{ct}},$
@@
$p_{k}(t)\leq\frac{c_{n}(m)}{(a-r_{k})\sqrt{ct}}\leq\delta.$
where $c_{n}(m)=\sqrt{3}\sigma_{n}(m)$ and obtain
$p=\max\left(\left\lfloor\frac{1}{2}+m-\frac{c_{n}(m)}{\delta\sqrt{ct}}\right\rfloor,0\right).$